Lower-of-Cost-or-Market Inventory

On the basis of the following data, determine the value of the inventory at the lower of cost or market. Assemble the data in the form illustrated in Exhibit 10.

Product

Inventory
Quantity

Cost Per
Unit


Market Value per Unit
(Net Realizable Value)

Class 1:   

Model A 19$115 $112

Model B 39  142   138

Model C 39294    305

Class 2:   

Model D 8  260    246

Model E 28 170   176

a. Determine the value of the inventory at the lower of cost or market applied to each item in the inventory.

Inventory at the Lower of Cost or Market




Product
Inventory
Quantity
Cost
per UnitMarket Value
per Unit
(Net Realizable Value)CostMarketLower of Cost or Market

Model A$$$$$

Model B

Model C

Model D

Model E

Total   $$$

b. Determine the value of the inventory at the lower of cost or market applied to each class of inventory.

Inventory at the
Lower of Cost
or Market




Product
Inventory
QuantityCost
per UnitMarket Value
per Unit
(Net Realizable Value)CostMarketLower of Cost or Market

Class 1:      

Model A$$$$

Model B

Model C

Subtotal   $$$

Class 2:      

Model D$$

Model E

Subtotal   $$

Total   $$$

c. Determine the value of the inventory at the lower of cost or market applied to total inventory.

Inventory at the Lower of Cost or Market




Product
Inventory
Quantity
Cost
per UnitMarket Value
per Unit
(Net Realizable Value)CostMarketLower of Cost or Market

Model A$$$$

Model B

Model C

Model D

Model E

Total   $$$

Check My Work
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Exhibit 10.
Answer a
Individual Items A B C=A*B D E=A*D F= Lower of C or E
Product Units Cost Per unit Cost NRV per unit NRV Lower of Cost or Market
Model A 19.00            115.00        2,185.00          112.00        2,128.00                  2,128.00
Model B 39.00            142.00        5,538.00          138.00        5,382.00                  5,382.00
Model C 39.00            294.00      11,466.00          305.00      11,895.00               11,466.00
Model D     8.00            260.00        2,080.00          246.00        1,968.00                  1,968.00
Model E 28.00            170.00        4,760.00          176.00        4,928.00                  4,760.00
Ending Inventory               25,704.00
Answer b
Class wise A B C=A*B D E=A*D F= Lower of C or E
Product Units Cost Per unit Cost NRV per unit NRV Lower of Cost or Market
Model A 19.00            115.00        2,185.00          112.00        2,128.00
Model B 39.00            142.00        5,538.00          138.00        5,382.00
Model C 39.00            294.00      11,466.00          305.00      11,895.00
Class 1     19,189.00     19,405.00               19,189.00
Model D     8.00            260.00        2,080.00          246.00        1,968.00
Model E 28.00            170.00        4,760.00          176.00        4,928.00
Class 2        6,840.00        6,896.00                 6,840.00
Ending Inventory               26,029.00
Answer c
Total Inventory A B C=A*B D E=A*D
Product Units Cost Per unit Cost NRV per unit NRV
Model A 19.00            115.00        2,185.00          112.00        2,128.00
Model B 39.00            142.00        5,538.00          138.00        5,382.00
Model C 39.00            294.00      11,466.00          305.00      11,895.00
Model D     8.00            260.00        2,080.00          246.00        1,968.00
Model E 28.00            170.00        4,760.00          176.00        4,928.00
Total Inventory     26,029.00     26,301.00
Ending Inventory- Lower of C or E= $ 26,029.